OpenStudy (anonymous):

an auditorium is 1/2 full. after 25 people in the audience left, the auditorium hall is 3/8 full. What is the seating capacity of the hall?

OpenStudy (dumbcow):

(1/2)A -25 = (3/8)A (1/2)A -(3/8)A = 25 (4/8)A - (3/8)A = 25 (1/8)A = 25 A = 25*8 = 200

OpenStudy (anonymous):

same solution with dumbcow (lol. what a name)